OpenStudy (anonymous):

please help with 6th grade math? I don't understand this at all!

OpenStudy (anonymous):

ok what can i do

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Evergreen bank offers 5% simple interest on deposits. Pot O' Gold bank offers 4% interest compounded annually. Patrick deposited $2000 into each bank to remain 3 years. A. Find the final balance at Evergreen Bank. Show All Your Work

OpenStudy (anonymous):

I have a b and c too i just wanna do on at a time

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

For simple interest, the formula is I=PRT/100 where P= Present value R= Interest rate T= Time ( no. of years) I= Interest So, for Evergreen Bank, the interest would be I=(2000×5×3)/100=$300 The final balance would be 2000+300=$2300

OpenStudy (anonymous):

For compound interest, the formula is, FV=P×(1+r)n here FV= Future value P= Present value r= Annual interest rate ( as a decimal) n= no. of years So, for Pot O' Gold bank, the final balance would be: FV=2000×(1+0.04)3=$2249.728

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

no prob you are welcome

OpenStudy (anonymous):

i know how to do the rest now thanks :D

OpenStudy (anonymous):

ok anything else just tag me and i will do my best to help answer